OpenStudy (anonymous):

simplify the radical expression √7 + 3 √7

jimthompson5910 (jim_thompson5910):

Let \[\Large x = \sqrt{7}\]

jimthompson5910 (jim_thompson5910):

So that means the expression turns into x + 3x Combine like terms to get 4x

jimthompson5910 (jim_thompson5910):

So that show us that \[\Large \sqrt{7} + 3\sqrt{7} = 4\sqrt{7}\]

jimthompson5910 (jim_thompson5910):

you can think of it as "I have 1 square root and then I have another 3 square roots...so I have 4 total square roots" the "square roots" have to be the same of course

OpenStudy (anonymous):

thank you!!! you're a life saver!

jimthompson5910 (jim_thompson5910):

you're welcome